2. A monthly record of glucose level of patients was recorded and found to have a mean of


98 mg/dL and a standard deviation of 10 mg/dL. If the data is normally distributed,


a) what percent of the patients has blood glucose level higher than 102 mg/dL?


b) what percent of the patients has blood glucose level between 120 and 125 mg/dL?


c) Assuming there are 1,500 patients tested, how many of the patients would have


glucose level above normal (glucose level ≥ 120 mg/dL)?

1
Expert's answer
2021-11-22T19:09:39-0500

a)


"P(X>102)=1-P(Z\leq\dfrac{102-98}{10})"

"=1-P(Z\leq0.4)\approx0.344578"


"34.4578\%"


b)


"P(120<X<125)=P(X<125)-P(X\leq120)"

"=P(Z<\dfrac{125-98}{10})-P(Z\leq\dfrac{120-98}{10})"

"=P(Z<2.7)-P(Z\leq2.2)"

"\approx0.9965330-0.9860966"

"\approx0.010436"

"1.0436\%"


c)


"P(X\geq120)=1-P(Z<\dfrac{120-98}{10})"

"=1-P(Z<2.2)\approx0.013903"

"0.013903(1500)=21"

21 patients.